OpenStudy (anguyennn):

How to evaluate √2/3?

OpenStudy (anonymous):

is it \[\frac{ \sqrt{2} }{ 3 }\]

OpenStudy (anguyennn):

yes

OpenStudy (jackthegreatest):

just think of it as \[\sqrt{2} \times \frac{ 1 }{ 3 }\]

OpenStudy (anonymous):

It's an irrational number, it cannot be evaluated, you can only approximate it, \[\sqrt{2} \approx 1.4142\] so we have \[\frac{\sqrt{2}}{3} \approx \frac{1.414}{3}=0.471333....\]

OpenStudy (anguyennn):

are you sure?

OpenStudy (anonymous):

Usually even upto 3 decimal digits is a good approximation if required

OpenStudy (anguyennn):

then what about 3/√3

OpenStudy (anguyennn):

how would you evaluate this?

OpenStudy (anguyennn):



OpenStudy (anonymous):

\[\frac{3}{\sqrt{3}}=\frac{\sqrt{3}\sqrt{3}}{\sqrt{3}}=\sqrt{3} \approx 1.732\]

OpenStudy (anonymous):

Notice that if you square 1.732, you get \[(1.732)^2=2.999824 \neq 3\]

OpenStudy (anonymous):

So it is only an approximation

OpenStudy (anguyennn):

but at the back of my answers it is showing √3 as the answer

OpenStudy (anonymous):

 \[\frac{ \sqrt{2} }{ 3 }=\frac{ 1.414 }{ 3 }=0.471 \approx\]